OpenStudy (anonymous):

3[2-4(2x-1)]=4x-10 3[2(-8x+4)]=4x-10 help? next step.

OpenStudy (lgbasallote):

multiply 2(-8x + 4)

OpenStudy (anonymous):

3[16x+8]=4x-10 ? correct right? then multiply the 3?

OpenStudy (lgbasallote):

no and yes

OpenStudy (anonymous):

-16**

OpenStudy (lgbasallote):

better

OpenStudy (lgbasallote):

now multiply 3

OpenStudy (anonymous):

check my math pls i got a fraction O.O

OpenStudy (anonymous):

-48x+24=4x-10 then you combine and change signs and i get -52x=-34?

OpenStudy (lgbasallote):

lol. scared of fractions?

OpenStudy (anonymous):

VERY!

OpenStudy (lgbasallote):

ohh..i see what went wrong

OpenStudy (lgbasallote):

\[3[2 - 4(2x-1)] = 4x - 10\] distribute -4 \[\implies 3[2 - 8x + 4] = 4x - 10\] you wrote it as 3[2(-8x+4)]

OpenStudy (lgbasallote):

so i thought it was multiplication...the next step should have been to combine like terms

OpenStudy (lgbasallote):

\[3[2 - 8x + 4] = 4x - 10\] \[\implies 3[6-8x] = 4x-10\] following so far?

OpenStudy (anonymous):

ohhhh okay i see what im doing wrong. I had it down like that as well. yes i am.

OpenStudy (lgbasallote):

good. so do you know what to do next?

OpenStudy (anonymous):

yes :D thank you!

OpenStudy (lgbasallote):

welcome